OpenStudy (anonymous):

What is the value of f(x) = 20 (2x) when x = 3

OpenStudy (♪chibiterasu):

Replace x with 3: f(3) = 20(2 • 3) f(3) = 20(6) f(3) = 120 Now solve for f: 3f/3 = 120/3 f = 40 Now check: 40(3) = 120 120 = 120